A mother is 30 years older than her daughter. Five years ago she was four times as old as her daughter. How old are they now?
Answer by ankor@dixie-net.com(15624)

Write an equation representing each statement
:
"A mother is 30 years older than her daughter."
M = D + 30
:
"Five years ago she was four times as old as her daughter."
(M-5) = 4(D-5)
Simplify:
M - 5 = 4D - 20
M = 4D - 20 + 5
M = 4D - 15
:
How old are they now?
4D - 15 = M
Substitute (D+30) for M in above equation, find D
4D - 15 = D + 30
4D - D = 30 + 15
3D = 45
D = 45/3
D = 15 yrs old is the daughter
M = 15 + 30 = 45 yrs old is the mom
;
Check solution using the statement:
"Five years ago she was four times as old as her daughter."
45 - 5 = 4(15-5)
:
Did I make this understandable to you?